# From Calculation to Adjudication: Examining LLM judges on Mathematical Reasoning Tasks

### v1

@@ 4.2 Performance per model combination
The comparative performance of model pairs offers insights into which model is better for the specific task or which combination of models yields the best results.
@@
Figure 2 illustrates the final performance $P(\Delta_{J}=T|A,B)$, indicating the probability of a judge choosing a correct answer given two models $A$ and $B$. The results are averaged over datasets and presented as an upper triangular matrix due to symmetry. If both models in a pair are the same, $A=B$, we employ the second response generated with temperature sampling to introduce variation. We report the performance of all models used as judges in the Appendix B in Table D.
@@ Results.
We observe that the best performance is achieved when both the candidate answers and the judge are the highest-performing model, Qwen 2 72B. An analysis of the first rows (cf. Figure 2) reveals a notable trend: The final performance declines when comparing the output of a strong model against a mediocre model (e.g., Llama3-8B) but then improves again when compared against the weakest model. This suggests that judging becomes more challenging when distinguishing between the correct answers of a strong model and the incorrect answers of a mediocre one, compared to discerning the outputs of a bad one.
@@ 4.3 Do judges elicit task improvement?
Figure 3: Amount of model pairs $(M_{A},M_{B})$ where the answers chosen by the judge achieve a higher task performance than the models individually (green). The blue bar only considers models where the judge is at least as good as the candidate models.
One use case for LLM judges is to improve task performance. A potential application is to train on answers chosen by the judge Yuan et al. (2024).
@@
Therefore, we test how often the performance of the answers chosen by the judge is better than the performance of the individual models. Formally, for all pairs of models $M_{A},M_{B}$ and datasets $D$, how often is the observed performance $P(\Delta_{J}=T|A,B,D)$ larger than $\max\{P(A=T|D),P(B=T|D)\}$? In Figure 3 the green bar tests all model pairs, and the blue bar only pairs where the judge is at least as good as the candidate models, i.e., $P(J|D)\geq\max\{P(A|D),P(B|D)\}$. The task performances of all models are given in the Appendix B in Table 9.
@@ Results.
We see that only Qwen 2 72B increases the performance reliably. However, it is easier for the judge to improve performance if it compares answers of less or equally good candidate models.
Figure 4: Percentage of model pairs $(M_{A},M_{B})$ where a judge picks a better model $M_{A}$ (meaning $P(A=T|D)>P(B=T|D)$), by selecting more answers of $M_{A}$ than from $M_{B}$.
4.4 Does the judge prefer the better model?
Another application of LLM judges is whether they can accurately identify which model performs better for a given task. This is crucial if we want to rank LLMs by their capabilities or if a practitioner wants to decide which model to deploy.
Setup.
To assess this, we evaluate the frequency with which a judge selects the superior model. For a candidate model pair $M_{A},M_{B}\in\mathbf{M}$, always assume they are ordered, such that $P(A=T|D)>P(B=T|D)$. Then, specifically, we determine the proportion for which the judge chooses $M_{A}$ more often than $M_{B}$, or formally, how often is $P(\Delta_{J}=M_{A}|A,B,D)>P(\Delta_{J}=M_{B}|A,B,D)$ for all candidate pairs and datasets.
Results.
The judges consistently perform well in the selection of the better model. Notably, we find that Qwen 2 72B can only not rank the pair Mistral 7B v0.1 and v0.3 on the MATH dataset. This issue appears minor, as both models exhibit similarly poor performance on the challenging MATH dataset (with accuracies of $6.13\%$ and $3.10\%$, respectively), meaning most judgments compare two wrong answers. Notably, already the worst judge, Mixtral 8x7B, performs well. In summary, we see that judges are more capable of aggregate-level rankings than instance-level rankings.
(a) Qwen2 72B
(b) LLama2 70B
(c) Yi 1.5 43B
(d) Mixtral 8x7B
Figure 5: Judges’ accuracy vs. performance gap between two candidate models $M_{A}$ and $M_{B}$. Each point represents a subset where $M_{A}$ is correct and $M_{B}$ is incorrect. The color reflects the size of these subsets.
5 Analysis of Subsets
We investigate properties that occur when we use subsets based on the correctness of models or agreement between models.
5.1 Do task performances correlate with judgments?
We consider the subset of highest practical relevance where one candidate model is correct, and one candidate model is incorrect. The goal is to investigate the relationship between candidate model task performance and judgment performance.
Setup.
For all model pairs $M_{A},M_{B}\in\mathbf{M},M_{A}\neq M_{B}$ we analyze subsets where $M_{A}$ is correct, and $M_{B}$ is incorrect. Note that we can always order $M_{A}$ and $M_{B}$ this way. Each plot in Figure 5 shows the relationship between judge performance, $P(\Delta_{J}=T|A=T,B=F)$ (Y-axis) and candidate model performance gap of $M_{A}$ and $M_{B}$, i.e., $P(A=T|D)-P(B=T|D)$ (X-axis). Examples of these subsets and their corresponding performances are in Appendix C in Table 10.
Results.
The analysis reveals a strong correlation (Pearson’s $r^{2}>0.69$) between candidate model performance gap and judgment accuracy. If the performance gap is negative, we consider subsets where larger models are incorrect. Judges favor answers from larger models even when they are incorrect on these subsets. We hypothesize that this bias arises because larger models exhibit a specific writing style, articulating their responses more convincingly, thereby misleading the judges. This finding aligns with previous research identifying self-bias Xu et al. (2024); Panickssery et al. (2024); Liu et al. (2024). However, our results indicate that this bias extends more broadly to the inherent quality of the underlying models on reasoning datasets. However, this is not necessarily a critical issue in practice, as the larger model tends to answer correctly more often (as indicated by the color of the points in Figure 5.
5.2 Does judgment quality depend on models’ agreement?
(a) Performance on all comparisons.
(b) Performance on comparisons with one correct and one incorrect answer.
Figure 6: Judge performance by agreement bucket. E.g., bucket $S_{3}$ (X-axis) means that all eight models gave together three different answers. Note that AQUA-RAT is multiple-choice with maximally six answers.
We are interested in whether the level of agreement among models, i.e., how many models give a different answer for a sample, impacts the performance on the respective subset.
Setup.
We define disagreement buckets $S_{j}$, where each bucket contains instances for which exactly $1\leq j\leq 8$ unique answers were given across all models. Formally, we set
$S_{j}=\bigcup_{D}\left\{i\in D\mid|\{M_{A}(i)\mid M_{A}\in\mathbf{M}\}|=j\right\}$
where $M_{A}(i)$ is the answer of model $M_{A}$ for instance $i$. We analyze the results in two contexts: all comparisons, including those where both answers are correct or incorrect (cf. Figure 6(a)), and only instances where exactly one answer is correct (cf. Figure 6(b)). We average the performances of all judges and all candidate pairs. Find per-judge plots in Appendix C in Figure 10.
Results.
Figure 6(a) shows that when all models agree (bucket $S_{1}$), the performance is nearly 100%, indicating unanimous agreement usually means correctness. As disagreement increases, performance expectedly decreases. Thus, model agreement is a proxy for sample difficulty. In 6(b), where a correct and an incorrect answer exists, performance remains relatively stable across disagreement buckets for datasets with free-form answers, such as GSM8K and MATH. However, for AQUA-RAT, performance degrades as disagreement rises.
[…]
↓ Features \  Condition → $A,B$ $A\neq B$$\{T,F\}=\{A,B\}$
(1) $P(J),P(A),P(B)$ 97.50 90.20 59.20
(2) $P(A=B)$, 76.00 54.90 49.90
$P(J=A|A\neq B)$
Table 2: Coeffictions of Determination ($R^{2}$, higher is better) for linear regression using the different feature sets as covariates (rows) and different target variables defined by the condition (columns). All values are significant ($p<0.001$) as per an Overall-F-Test.
6 Prediction of Judgements
We investigate whether predicting the judgments’ outcomes is feasible. Firstly, we aim to predict performance statistics. Secondly, we aim to predict individual judgments.
6.1 Can we predict judgment performance?
On the subset where exactly one answer is correct, we found a strong correlation between judgment performance and candidate task performances. This hints at regularities within the judging process, thus we aim to predict judge performance using model statistics.
Setup.
We fit six different linear regression models using the judgment performances as the target variables $Y$, including all variations of judges, model pairs $M_{A},M_{B}\in\mathbf{M}$, and datasets $D$. Regarding the covariates $\mathbf{X}$ in the model, we distinguish between two setups: In Case (1), we solely use the task performances $P(X|D),X\in\{J,A,B\}$ of judge and candidate models, to predict judgment performance. In Case (2), we utilize statistics available without knowledge of the ground truth. The features for this case are the probability of agreement between the candidate models $P(A=B|D)$ and the probability of model $M_{A}$ being chosen. Since we are not specifically interested in the individual features’ effects, but rather in their ability to explain the variation of judgment performance, we rely on the coefficient of determination, $R^{2}$, for evaluation (Fahrmeir et al., 2013, , see Appendix E).
Results.
The results are shown in Table 2 (excluding data sets from the probability formulas for simplicity). We observe that the performance-related features of the models can almost perfectly explain the variation in final judgment performance ($R^{2}=97.50\%$), also when conditioning only on the subset of differing answers ($R^{2}=90.20\%$). Logically, $P(A)$ and $P(B)$, i.e., $P(A|D),P(B|D)$ respectively, have significant [^3] We test statistical significance using an Overall-F-Test for each fitted model. Further details are in Appendix E. explanatory power for judgment performance, as they encompass all correct answers. In Case (2), we still observe a relatively high $R^{2}$ value, indicating that the features can explain 50% of the target’s variance.
[…]
(1) TF-IDF + RF 60.78 61.37 60.77 58.69
(2) RoBERTa 68.14 66.49 67.03 63.91
Table 3: Accuracy of predicting LLM judges’ decisions using Random Forest (RF) and RoBERTa classifiers.
6.2 Can we predict which individual judgments?
We hypothesize that judgments are biased towards larger or better models because they incorporate linguistic cues or writing style into their judgments rather than purely relying on reasoning assessment. Therefore, we train a classifier to understand whether we can predict individual judgments.
Setup.
We separate all comparisons made per judge into training, validation, and test splits and train two classifiers. The test accuracy is reported in Table 3. The first model utilizes TF-IDF vectorization. We create two independent vectorizers for both answers. The resulting features are concatenated. A RandomForest classifier Breiman (2001) is then trained on these combined features. The second model is a RoBERTa model Liu et al. (2020) trained on the full prompt presented to the judge. Refer to Appendix D for the training details of both models.
Results.
The random forest model achieves an accuracy of approximately 60%, demonstrating performance above random chance. This suggests that specific keywords or phrases influence judges. The RoBERTa model surpassed this, reaching nearly 70% accuracy. Taken together, these results suggest that judge decision-making is a multi-faceted process. While specific linguistic cues appear to hold influence, a substantial portion of the decision-making process seems to be based on other contextual factors or broader reasoning.
7 Perturbation of Results
Swapped Masked
= $\neq$ Refused = $\neq$ Refused
Qwen2-72B 75.75 12.20 12.05 56.40 26.06 17.53
Llama-3-70B 78.81 13.72 7.47 63.44 29.22 7.33
Yi-1.5-34B 74.80 14.47 10.73 44.89 37.31 17.80
Mixtral-8x7B 71.19 20.69 8.12 60.34 29.26 10.40
Table 4: Analysis of judgments where results in candidate answers were either swapped or numbers masked. We report how many judgments stay the same (=)(=)( = ), different ()($\neq$)( ≠ ), or where judges refused to follow the output format (Refused).
We aim to gain a deeper understanding of the extent to which writing style affects the final judgment. Therefore, we create an experiment perturbing the candidate answers and examine whether this changes the judgment.
Setup.
We examine two perturbtations: Swap and Mask. In the Swap experiment, we swap the final answer from model $M_{A}$ with that of model $M_{B}$, while keeping their CoT reasoning unchanged. In the Mask experiment, we anonymize all numbers in both the CoT reasoning and the final answer by replacing them with “X”. [^4] In preliminary runs, we observed that masking caused significant confusion for the judge models. To address this, we adapt the judgment prompt in this setting to include the instruction: “Only analyze the reasoning! All numbers have been replaced with ’X’ to help you focus on the reasoning.” Table 4 shows the frequency with which the judge selects the same answer ($=$), a different answer ($\neq$), or fails/refuses to follow the output format and make a decision (Refused). Refer to Appendix F for specific examples.
Results.
We observe that the new judgments in more than half the cases agree with the original judgment. In the Swap experiment, they even agree on average by 75% of the cases. We deduce that the judge is largely unaffected by the artificially introduced noise and heavily bases its decision on the writing style. Interestingly, in a substantial amount of samples (up to $17\%$) the judge refuses to make a judgment. On a positive note, manual inspection revealed that the model often realizes that the original answers were perturbed.
8 Discussion
Style and Quality.
Our experiments suggest a relation between judgment and candidate task performance (cf. Section 5) and a relation between judgment and writing style (cf. Sec. 6 and 7). We hypothesize these two are interconnected and facets of the same underlying bias. When models become better, e.g., by being trained on larger amounts of data, their ability to write convincingly increases. Conversely, when an LLM demonstrates an increased ability to write convincingly, it likely acquires a more nuanced grasp of what humans perceive as compelling. This enhanced understanding likely also extends to task performance.
Generalizability of approach.
Our in-depth analysis utilizes Formula (1) to segment judgment data based on correctness criteria, allowing for targeted investigation of specific subsets. This approach is generalizable and transferable to other NLP tasks, such as summarization. By incorporating discrete signals such as text topics, a similar derivation of the judgment probability is possible.

### v2

@@ 4.2 Performance per model combination
Each model has unique strengths and weaknesses and often answers different questions correctly. In this section, we analyze the judgment performance per model pair to gain a better understanding of the impact of candidate model combinations on judgment performance.
@@
Figure 3 illustrates the judgement performance $S^{J}_{A,B}$ across model pairs $(A,B)$, indicating the probability of a correct judgement. The results are averaged over datasets and presented as an upper triangular matrix due to symmetry (we always present the answers in both possible orders and average performance). We report the performance of all models used as judges in the Appendix B in Table 9.
@@ Results.
The highest performance is achieved when two answers of Qwen 2.5 72B are compared which is the highest performing model (see task performance in Appendix B.1) In general, we observe that it is easier for the judge to make a correct judgment if candidate models are of higher quality. This seems intuitive because such models likely structure and present their reasoning well, allowing a judge to compare solutions more easily. Figure 2 gives an additional explanation. It shows that judges very reliably detect whether both answers are correct. When both models are capable, it is more likely that both give a correct answer, which makes it easier for LLM judges to classify correctly.
(a) LLama 3.1 70B
(b) Qwen 2.5 72B
(c) Qwen 2.5 14B
(d) Gemma 2 27B
Figure 3: Judgment Performance $S^{J}_{A,B}$ of LLM judges on model pairs, averaged across datasets.
Interestingly, the judgment performances of Qwen 2.5 72B, Qwen 2.5 14B, and LLama 3.1 70B are very similar across pairs. The former possibly agree on a lot because of the similarity of training data and knowledge distillation Hinton et al. (2014). The largest performance difference is that LLama 3.1 70B performs 10% better when Qwen 2.5 72B and Gemma 2 2B are compared.
These results show that there is a relationship between the task performance of a candidate model and judgment performance. The following section will provide further analysis.
5 Population-level Analysis: Judgements and Model Quality
In this section, we investigate the relationship between LLM judgments and candidate LLM quality. First, we provide a statistical analysis where we use LLM task performance to explain the variance in LLM judgment performance. Further, we focus on the subsets where the candidate models make exactly one correct and one incorrect prediction. We observe a strong statistical relationship between the difference in candidate task performances and judgment performances.
5.1 Can we explain Judgement Performance using Task Performance?
A good indicator of the competence of a model on a specific dataset is its task performance. Clearly, there is a relationship between the quality of the involved models and the made judgments. We investigate the relationship between task performances (of candidate and judge models) and judgment performance.
Llama 3.1 70B Qwen 2.5 72B Qwen 2.5 14B Gemma 2 27B
$R^{2}$ 0.89 0.87 0.85 0.93
($p$-value) (0.0) (0.0) (0.0) (0.0)
Table 2: $R^{2}$ values for the regression models per judge (first row) and corresponding $p$-values of the Overall F-Test (second row). All $R^{2}$ values are statistically significant on the 5% level.
Setup.
We fit multiple different linear regression models using the judgment performances as the target variables $Y$, including all variations of judges, model pairs $(A,B)$, and datasets $D$. Regarding the covariates $\mathbf{X}$ in the model, we solely use the task performances $S_{X},X\in\{J,A,B\}$ of judge and candidate models, to predict judgment performance. Since we are not specifically interested in the individual features’ effects, but rather in their ability to explain the variation of judgment performance, we rely on the coefficient of determination, $R^{2}$, for evaluation (Fahrmeir et al., 2013, see Appendix D).
Results.
The results are shown in Table 2 (excluding data sets from the probability formulas for simplicity). We observe that the performance-related features of the models can explain the variation in the judgment performance (all $R^{2}$ values greater or equal than 0.85), very well. Logically, $S_{A}$ and $S_{B}$, have significant [^3] We test statistical significance using an Overall-F-Test for each fitted model. Further details are in Appendix D. explanatory power for judgment performance, as they encompass all correct answers.
5.2 Are LLM judges biased towards LLMs of higher quality?
To get a better understanding of whether there is a bias of LLM judges towards LLMs of higher quality, we investigate the subset where one candidate answer is correct and the other candidate answer is incorrect. This subset is of the highest practical relevance. The goal is to investigate the relationship between the task performances of the candidate models and the judge’s performance.
Setup.
For all model pairs $(A,B),A\neq B$ we analyze subsets where $A$’s solutions are correct, and $B$’s solutions are incorrect, and call it 1-correct. Note that we can always order $A$ and $B$ this way. Each plot in Figure 4 shows the relationship between judge performance on the 1-correct subset (Y-axis) and candidate model performance gap of $A$ and $B$, i.e., $S_{A}-S_{B}$ (X-axis). The color of the points indicate the size of the particular subset of samples. Examples of these subsets and their corresponding performances are in Appendix C.1.
Results.
The analysis reveals a strong correlation (Pearson’s $r^{2}>0.78$) between judgement performance and candidate model performance gap. For the rest of this section, we call the model of higher performance on a dataset the more competent model. I.e., if the performance gap is larger than 00, the model giving the correct answer ($A$) is the more competent model. If the correct model is the more competent model, the judgment performance on the subset is higher, e.g., for LLama 3.1 70B, sometimes approaching 100%. If the performance gap is more positive, it is easier to choose the correct answer. On the other hand, if the less competent model gives the correct answer, judgment performance is low, often lower than $20\%$.
(a) Llama 3.1 70B
(b) Qwen2.5 72B
(c) Qwen 2.5 14B
(d) Gemma 2 27B
Figure 4: Judges’ accuracy vs. performance gap between two candidate models $A$ and $B$. Each point represents a subset where $A$ is correct, and $B$ is incorrect. The color reflects the size of these subsets.
We infer that LLM judges are biased towards models of higher task performance. This finding aligns with previous research identifying self-bias Xu et al. (2024); Panickssery et al. (2024); Liu et al. (2024), as judge LLMs are typically of higher quality than the judged models. We hypothesize that this bias arises because more competent models articulate their responses more convincingly and exhibit a specific writing style, thereby misleading the judges.
However, models of higher task performance typically answer correctly more often (as indicated by the color of the points in Figure 4.
6 Sample-level analysis: Judgments and Stylistic Patterns
In Section 5, we found that the quality of a candidate LLM (as indicated by the task performance) correlates with the made judgment. We hypothesize that models of higher quality exhibit a particular style of expressing themselves and judges partially base their judgment on the incorporated textual cues. Motivated by recent work in machine-generated text detection which finds that LLMs often exhibit certain styles Wu and Aji (2025) or patterns Shaib et al. (2024), we aim to gain a deeper understanding of whether shallow or even content-independent patterns affect the final judgment.
Setup.
We separate all judgments each judge made into training and test splits and train two classifiers. The test accuracy is reported in Table 3. We use two types of features. First, we use TF-IDF embeddings. Secondly, we use N-Grams of part-of-speech (POS) tags, motivated by Shaib et al. (2024) who show and investigate the distinct occurrence of such in LLM-generated text. Given two candidate answers, we create two independent feature sets and concatenate those. Then a logistic regression and a RandomForest classifier Breiman (2001) are trained on these concatenated features. Find more information in Appendix E.
Results.
We observe that the models achieve a performance between approximately 70% and 75%. This indicates that structural information (POS tags) and word choice (TF-IDF) are important factors in understanding the patterns behind the behavior of LLM judges. The ground truth judgment distribution is shown in Appendix E.
Features Model Llama 3.1 70B Qwen 2.5 72B Qwen 2.5 14B Gemma 2 27B
POS LR 72.79 69.66 72.33 70.19
RF 71.71 69.77 71.89 69.18
TF-IDF LR 75.75 73.65 75.12 72.27
RF 75.65 71.05 75.79 70.58
Table 3: Accuracy of predicting LLM judges’ decisions using Logistic Regression (LR) and Random Forest (RF) classifiers based on N-Grams of either POS tags or TF-IDF features.
Nevertheless, these results suggest that decision-making is a multi-faceted process. While specific shallow cues hold influence, a substantial portion of the decision-making process (25%-30%) can not be predicted this way and is based on other contextual factors which could include reasoning or noise.
7 Usage recommendations
Lastly, we aim to give some usage recommendations. We start by analyzing two applied questions, namely, whether LLM judges can identify models of higher task performance and whether LLMs should be used to improve task performance. In the end, we discuss those results, connecting them to the overall insights of this paper.
7.1 Do judges identify better models?
An essential application of LLM judges is whether they can accurately identify which model performs better for a given task. This is crucial if we want to rank LLMs by their capabilities or if a practitioner wants to decide which model to deploy.
Setup.
We evaluate which model a judge perceives as better by measuring the frequency of how often a judge selects the answer of a specific model. Formally, let $(A,B)$ be a candidate model pair where we assume that $A$ has higher task performance, i.e. $S_{A}>S_{B}$. If the judge chooses $A$ more often, we say a judge correctly determines $A$ to be better than $B$. For this analysis, we determine the proportion of model pairs $(A,B)$ for which the judge chooses $A$ over $B$ for all pairs $(A,B),S_{A}>S_{B}$ as shown in Figure 5.
Figure 5: Percentage of model pairs $(A,B)$ where a judge picks a better model $A$ (meaning $S_{A}>S_{B}$), by selecting more answers of $A$ than from $B$.
Results.
We observe that all tested large models consistently select the more competent model, i.e., the model with higher task performance. Also, the small models with $7$-$9$B parameters choose the correct model in over $90$% of the cases. In general, it seems to be the hardest on the AQUA-RAT dataset. This is also the hardest dataset in Case (4), in Table 1, where exactly one answer is correct. Note that the bias found in Section 5.2 is not necessarily problematic for this specific use case, because a bias towards the more competent model supports a correct outcome of this experiment.
@@ 7.2 Do judges elicit task improvement?
Another interesting question of practical relevance is whether it makes sense to use LLM judges to improve task performance. One use case is the application of LLM judges in agentic systems where LLM judges might serve as a dedicated unit in a system. Another use case is the subsequent usage of the answers chosen by the judge for self-training Yuan et al. (2024).
@@
We separate the analysis into two questions. In Case (1), we evaluate whether the answers chosen by the judge result in a better performance than the individual models. Formally, for all pairs of models $(A,B)$, we plot the difference of performance of chosen answers, $C^{J}_{A,B}$ and maximal single candidate model performance $\max\{S_{A},S_{B}\}$ in blue in a bar chart in Figure 6. Secondly, in Case (2), we test whether it makes more sense to use the judge model to generate a candidate answer $J$ and then take the majority vote across all three answers. Therefore we plot the performance difference of $C^{J}_{A,B}-\text{MV}(A,B,J)$ in orange in a bar chart, where $\text{MV}(A,B,J)$ is the performance of the majority vote across all three answers.
@@ Results.
In general, we observe that the performance differences are almost following a normal distribution. In Case (1), the distribution has a mean value (dashed line) slightly larger than 0 for LLama 3.1 70B (0.3) and Qwen 2.5 14B (0.9). That means that, on average, the answers chosen by the judge result in slightly increased performance, e.g., an increase from 40% accuracy to 40.9% accuracy. In Case (2), the mean value is never larger than 0, meaning that the majority vote is more likely to be better than the answer chosen by the judge. Especially for Qwen 2.5 14B and Qwen 2 72B, it is more viable to use the majority voting strategy.
(a) LLama3.1 70B Full
(b) Qwen 2.5 72B Full
(c) Qwen 2.5 14B Full
(d) Gemma 2 27B
Figure 6: The Y-axis describes the number of model pairs $A,B$ where the answers chosen by the judge achieve a higher task performance than the performances of the individual models (blue) or than the majority vote (MV) of answers of $A,B$ and $J$ as candidate answer generator (red). The X-axis describes the performance difference. A value of, e.g., $x=0.05$ means the answers chosen by the judge result in a 5% (absolute) performance increase.
7.3 Discussion
Our analysis of LLM judges on mathematical reasoning tasks reveals several insights for practitioners, which we discuss in the following. We separate our discussion into the sample level, i.e., the interpretation of a single prediction, and aggregate level, i.e., the interpretation of a set of predictions.
Sample level.
In Table 1, we find that LLM judges often achieve a strong judgment performance ($S^{J}_{A,B}$ $>80\%$ accuracy) across tasks. While this is a solid classification performance, it means that the prediction is wrong in 20% of the cases which limits practical applicability. In Section 4, we observe that LLM judges demonstrate high precision when identifying correct answers from both models. This might be valuable for filtering samples and curating training data or, e.g., self-training. Nevertheless, one has to be careful how to use these because correctly judged samples are biased towards simple samples. In summary, we do not recommend fully relying on individual LLM judgments, especially not in high-stakes domains such as legal or health care.
Aggregate level.
As shown in Figure 5, we find that LLM judges are consistently able to select or rank models by their task performance. This is supported by Section 5.1 where we show that a simple linear model can explain a high share of the variance in judgment performance, given individual task performances, suggesting that the performance difference of two candidate models is linearly linked to the judgment outcome.
In summary, our results suggest that LLM judges are more effective and consistent at aggregate-level comparisons than instance-level judgments, for example when ranking or selecting which LLM is better for a particular task when no ground truth data is available.